OpenStudy (anonymous):

Point P is chosen at random in a circle. If a square is inscribed in the circle, what is the probability that P lies outside the circle

OpenStudy (unklerhaukus):

zero. the point is chosen inside the circle, which means it is not outside the circle

OpenStudy (anonymous):

thanks!

OpenStudy (unklerhaukus):

it that really the question? i was kinda hoping you wrote it wrong

OpenStudy (anonymous):

wait. it says the answers taht are possible for this question are A.) 1-1/2pie B.) 1-2/pie C.) 1-Pie/2 or D.) 1-1/4pie

OpenStudy (anonymous):

Is it possible to draw a diagram

OpenStudy (anonymous):

I messed up the question. that last word should say square not circle

OpenStudy (anonymous):

Point P is chosen at random in a circle. If a square is inscribed in the circle, what is the probability that P lies outside the (square)-is this correct

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Were you given units-dimensions

OpenStudy (anonymous):

no:/

OpenStudy (unklerhaukus):

Something like this  So to solve this problem you need to find the ratio of the areas

OpenStudy (unklerhaukus):

The diameter of the circle is the same as the diagonal on the square right?

OpenStudy (anonymous):

Im so lost...

OpenStudy (anonymous):

but UnkleRhaukus the square could be any size

OpenStudy (unklerhaukus):

The area of the circle is\[A_c=\pi r^2\]or\[A_c=\pi(d/2)^2\] The area of the square is \[A_s=s^2\]

OpenStudy (unklerhaukus):

and the diagonal on the square is\[d=\sqrt{s^2+s^2}\]

OpenStudy (unklerhaukus):

The area of the circle \[A_c=\pi \left( \frac{d}{2} \right)^2=\pi\left(\frac{\sqrt{2s^2}}{2}\right)^2=\pi \frac{s^2}{2}\]

OpenStudy (unklerhaukus):

So we have the formula for the area of the square and the area of the circle, both as functions of the same variable 's'. To find the portion of the circle covered by the square simple divide; A_s / A_c

OpenStudy (unklerhaukus):

the probability that the point P is not in this area but in the circle is \[1-\frac{A_s}{A_c}\] and you should get one of the options A,B,C,or D. i wont say which it is but i have one of those as my final answer. If i haven't explained this well let me know.

OpenStudy (unklerhaukus):